# Pseudo-Physics-Informed Neural Operators: Enhancing Operator Learning from Limited Data

### v1

@@ Abstract
Neural operators have shown great potential in surrogate modeling. However, training a well-performing neural operator typically requires a substantial amount of data, which can pose a major challenge in complex applications. In such scenarios, detailed physical knowledge can be unavailable or difficult to obtain, and collecting extensive data is often prohibitively expensive. To mitigate this challenge, we propose the Pseudo Physics-Informed Neural Operator (PPI-NO) framework. PPI-NO constructs a surrogate physics system for the target system using partial differential equations (PDEs) derived from simple, rudimentary physics principles, such as basic differential operators. This surrogate system is coupled with a neural operator model, using an alternating update and learning process to iteratively enhance the model’s predictive power. While the physics derived via PPI-NO may not mirror the ground-truth underlying physical laws — hence the term “pseudo physics” — this approach significantly improves the accuracy of standard operator learning models in data-scarce scenarios, which is evidenced by extensive evaluations across five benchmark tasks and a fatigue modeling application.
Machine Learning, ICML
@@ 1 Introduction
Operator learning, an important area for data-driven surrogate modeling, has made significant strides with the emergence of neural operators, which leverage the expressive power of neural networks. Notable examples include Fourier Neural Operators (FNO) (Li et al., 2020b, ), Deep Operator Net (DONet) (Lu et al.,, 2021) and other frameworks such as (Cao,, 2021; Hao et al.,, 2023). FNO employs Fourier transform for global convolution and function transformation, while DONet introduces two sub-networks, the branch net and trunk net, to extract representations from the functional space and query locations, respectively, enabling predictions akin to attention mechanisms (Vaswani et al.,, 2017).
For trading for model capacity and performance, neural operators typically require a substantial amount of training data to perform optimally. This demand poses challenges, particularly in complex problems, where training data is limited and costly to acquire. In response, the field of physics-informed machine learning, including physics-informed neural networks (PINN) (Raissi et al.,, 2019), has shown promise by incorporating physical laws as soft constraints during training. This approach serves as a regularization technique, embedding a fundamental understanding of physics into the model to lessen its reliance on extensive data. Building on this idea, the concept of physics-informed neural operators (PINO) has emerged, which integrates physical laws as soft constraints to enhance operator learning while reducing data quantity. It has been used in (Wang et al.,, 2021; Li et al.,, 2021) for FNO and DONet training.
@@
  • The effectiveness of PPI-NO is validated through extensive evaluations on five commonly used benchmark operator learning tasks in literature (Li et al., 2020b, ; Lu et al.,, 2022), including Darcy flow, nonlinear diffusion, Eikonal, Poisson and advection equations, as well as one application in fatigue modeling in fracture mechanics, where the ground-truth holistic PDE system is unknown.
@@ 2 Background
where ${\bf x}$ is a compact notation for the spatial and temporal coordinates, $\Omega$ is the spatial domain, $[0,\infty)$ is the temporal domain, $\mathcal{N}$ is a nonlinear differential operator, $u(\mathbf{x})$ is the solution function, and $f(\mathbf{x})$ is the source term. We aim to learn the solution operator of the PDE system, ${\psi}:\mathbb{F}\rightarrow\mathbb{U}$ where $\mathbb{F}$ and $\mathbb{U}$ are two functional spaces, using a training dataset $\mathcal{D}=\{({\bf f}_{n},{\bf u}_{n})\}_{n=1}^{N}$, which includes different instances of $u(\cdot)$ and $f(\cdot)$ sampled/discretized at a set of locations. Once the model is trained, it can be used to directly predict the solution function $u$ for new instances of the input $f$, offering a much more efficient alternative to running numerical solvers from scratch. However, the training dataset still needs to be generated offline using numerical solvers.
[…]
@@
Deep Operator Network (DONet) (Lu et al.,, 2021) is another prominent work in operator learning. The architecture of a DONet is structured into two components: a branch net and a trunk net, learning representations for the input functions and querying locations, respectively. Consider an input function $f({\bf x})\in\mathbb{F}$ sampled at $m$ locations $\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{m}\}$ and an output function $u\in\mathbb{U}$. The branch net receives the values $[f({\bf x}_{1}),f({\bf x}_{2}),\cdots,f({\bf x}_{m})]$ and outputs a feature representation $[b_{1},b_{2},\cdots,b_{p}]^{\top}\in\mathbb{R}^{p}$. Concurrently, the trunk network processes a querying location ${\bf x}$ and outputs another feature vector $[t_{1},t_{2},\cdots,t_{p}]^{\top}\in\mathbb{R}^{p}$. The output function value at ${\bf x}$ is predicted as a sum of products of the corresponding elements from the branch and trunk nets, $\mathcal{G}[f]({\bf x})\approx\sum_{k=1}^{p}b_{k}t_{k}$, where $\mathcal{G}$ is the learned operator mapping $f$ to $u$.
@@
Physics-Informed Neural Operator (PINO) (Wang et al.,, 2021; Li et al.,, 2021) has recently emerged as a promising approach to address the data scarcity issue in operator learning. PINO embeds physical laws — typically governing equations — into the learning process. The incorporation of physics not only enhances the model’s adherence to ground-truth phenomena but also reduces its dependency on extensive training data. Mathematically, the integration of physics into the learning process can be viewed as adding a regularization term in the loss function. Let $\mathcal{L}_{\text{data}}$ represent the standard data-fitting loss term (e.g., the mean squared error between the predicted and actual outputs), and the physics-informed term $\mathcal{L}_{\text{physics}}$ be the residual of the governing PDEs evaluated at the neural network’s outputs. The loss function $\mathcal{L}$ for a PINO model is expressed as
@@ 3 Methodology
In the absence of physics knowledge (i.e., PDE system (1)), it is impossible to construct the physics loss term as in PINO. To address this challenge, we propose a “pseudo” physics-informed operator learning framework that extracts useful physics representation from data so as to enhance operator learning. This framework is motivated by relatively complex applications, where data is often costly and/or limited while the underlying physics is hard to fully understand.
@@ 3.1 Pseudo Physics System Learning
As the first step, we propose a novel approach to learn the physics system using scarce training data. Our key observation is that, although the mapping from $f$ to $u$ can be intricate and may necessitate information across the entire domain (e.g., in linear PDEs, $u$ is an integration of the Green’s function multiplied with $f$ over the domain), the underlying PDE system (1) simplifies to a local combination of $u$ and its derivatives. We therefore design a neural network $\phi$ to approximate the general form of $\mathcal{N}$,
@@
The inherent local combinatorial nature of the PDE representation decouples the values of $u$ and its derivatives across different sampling locations, thereby significantly reducing the learning complexity and the amount of training data required. For instance, consider sampling the input function $f$ and output function $u$ on a $128\times 128$ grid. A single pair of discretized input and output functions, denoted as $({\bf f},{\bf u})$, is typically insufficient for a neural operator to effectively learn the mapping $f\rightarrow u$. However, this pair can be decomposed into $128\times 128=16,384$ training data points across various (spatial and temporal) locations, which can be used to train $\phi$ as outlined in (3). This decomposition provides rich information that reveals the local relationships between those derivatives. Therefore, even with a small number of $({\bf f},{\bf u})$ pairs, we hypothesize that the learning of the PDE system $\mathcal{N}$ through our formulation in (3) can still yield high accuracy in predicting $f(\cdot)$ as in (1).
@@
We use finite difference to obtain the derivatives of each $u_{n}$, namely, $S_{k}(u_{n})$ ($1\leq k\leq Q$), and then pass these inputs to the network $\phi$ to compute the prediction. Since finite difference introduces numerical errors, the computed derivatives are not exact. To compensate for such information loss, we propose incorporating a convolution layer to integrate and leverage the neighborhood information (see Fig. 1 top). Specifically, let $\hat{S}_{k}(u_{n})$ represents the finite difference approximation of $S_{k}(u_{n})$ ($1\leq k\leq Q)$; for $k=0$, we define $\hat{S}_{k}(u_{n})=u_{n}$. Each $\hat{S}_{k}(u_{n})$ is treated as an input channel. After applying the convolution, the output at each sampling location ${\bf x}_{j}$ for channel $k$ is given by
@@
where $\text{nei}(\cdot)$ denotes the neighborhood sampling locations defined in the convolution filter, $\Delta_{ij}$ is the relative distance between ${\bf x}_{i}$ and ${\bf x}_{j}$, and $w_{c}(\Delta_{ij})$ is the corresponding filter weight. We view $\overline{S}_{k}({\bf x}_{j})$ as an interpolation, providing a new approximation of $S_{k}({\bf x}_{j})$ by incorporating all neighborhood values of $u_{n}({\bf x}_{j})$ and their finite difference derivative approximations. The convolution results are then passed into subsequent layers. The interpolation (filter) weights are jointly learned from data. This design allows us to integrate additional information into the inputs to facilitate the learning of $\phi$. Empirical results confirm that incorporating this convolution layer improves the prediction accuracy; refer to the ablation study in Section 5.1 and Appendix Table 6(a). Next, at each sampling location ${\bf x}_{j}$, we employ fully connected layers (i.e., MLP) to combine all $\{\overline{S}_{k}({\bf x}_{j})\}_{k=0}^{Q}$ in an arbitrarily flexible way to predict $f({\bf x}_{j})$; see Fig. 1 bottom. In this way, the learned neural network mapping $\phi:u\rightarrow f$, although black-box in nature, is expressive enough to encapsulate valuable physics knowledge inherent in the data employed for operator learning.
Our method can be readily adapted to other numerical approaches for derivative approximation. For instance, when functions are irregularly sampled, smooth function estimators such as kernel interpolation (Long et al.,, 2024), the RBF-FD method (Fornberg et al.,, 2013; Fornberg and Flyer,, 2015), or Bayesian B-splines (Sun et al.,, 2022) can be used to estimate gradient information directly from data. These estimated gradients are then provided as inputs to our PDE neural network $\phi$ for further training.
@@ 3.2 Coupling Neural Operator with Pseudo Physics
Next, we leverage the pseudo physics laws embedded in the learned mapping $\phi:u\rightarrow f$ to enhance neural operator learning. Specifically, we use $\phi$ to reconstruct $f$ from the $u$ predicted by the neural operator. In this way, our approach goes beyond relying solely on the training data; it uses the physics learned in the previous step to incorporate a reconstruction error into the learning of the neural operator parameters; see Fig. 2 for an illustration.
@@
where $N^{\prime}$ is the number of input function samples.
@@
2:  Train a preliminary psuedo physics network $\phi$ with loss ${\mathcal{L}}_{\phi}$ in (3).
@@
5:     Fix pseudo physics network $\phi$, fine tune NO $\psi$ with the loss (5).
6:     Fix NO $\psi$, fine tune the pseudo physics network $\phi$ with the loss (5).
7:  until Maximum iterations are done or convergence
This methodology mirrors human experts’ approach to physics system modeling, where a limited amount of data is available and used to investigate the physics laws inspired by simple differential operations, and then these laws are utilized to generalize the system for further applications, e.g., forward simulation and solving inverse problems. In our method, the reconstruction loss term augments the operator learning with additional information, leading to potential improvement upon only training with limited data.

### v2

@@ Abstract
Neural operators have shown great potential in surrogate modeling. However, training a well-performing neural operator typically requires a substantial amount of data, which can pose a major challenge in complex applications. In such scenarios, detailed physical knowledge can be unavailable or difficult to obtain, and collecting extensive data is often prohibitively expensive. To mitigate this challenge, we propose the Pseudo Physics-Informed Neural Operator (PPI-NO) framework. PPI-NO constructs a surrogate physics system for the target system using partial differential equations (PDEs) derived from simple, rudimentary physics principles, such as basic differential operators. This surrogate system is coupled with a neural operator model, using an alternating update and learning process to iteratively enhance the model’s predictive power. While the physics derived via PPI-NO may not mirror the ground-truth underlying physical laws — hence the term “pseudo physics” — this approach significantly improves the accuracy of standard operator learning models in data-scarce scenarios, which is evidenced by extensive evaluations across five benchmark tasks and a fatigue modeling application. Our implementation is released at https://github.com/BayesianAIGroup/PPI_NO
@@ 1 Introduction
Operator learning, an important area for data-driven surrogate modeling, has made significant strides with the emergence of neural operators, which leverage the expressive power of neural networks. Notable examples include Fourier Neural Operators (FNO) (Li et al., 2020b), Deep Operator Net (DONet) (Lu et al., 2021) and other frameworks such as (Cao, 2021; Hao et al., 2023). FNO employs Fourier transform for global convolution and function transformation, while DONet introduces two sub-networks, the branch net and trunk net, to extract representations from the functional space and query locations, respectively, enabling predictions akin to attention mechanisms (Vaswani et al., 2017).
For trading for model capacity and performance, neural operators typically require a substantial amount of training data to perform optimally. This demand poses challenges, particularly in complex problems, where training data is limited and costly to acquire. In response, the field of physics-informed machine learning, including physics-informed neural networks (PINN) (Raissi et al., 2019), has shown promise by incorporating physical laws as soft constraints during training. This approach serves as a regularization technique, embedding a fundamental understanding of physics into the model to lessen its reliance on extensive data. Building on this idea, the concept of physics-informed neural operators (PINO) has emerged, which integrates physical laws as soft constraints to enhance operator learning while reducing data quantity. It has been used in (Wang et al., 2021; Li et al., 2021) for FNO and DONet training.
@@
  • The effectiveness of PPI-NO is validated through extensive evaluations on five commonly used benchmark operator learning tasks in literature (Li et al., 2020b; Lu et al., 2022), including Darcy flow, nonlinear diffusion, Eikonal, Poisson and advection equations, as well as one application in fatigue modeling in fracture mechanics, where the ground-truth holistic PDE system is unknown.
@@ 2 Background
where $\bf x$} is a compact notation for the spatial and temporal coordinates, $\Omega$ is the spatial domain, $0,\infty)$ is the temporal domain, $\mathcal{N}$ is a nonlinear differential operator, $u(\mathbf{x})$ is the solution function, and $f(\mathbf{x})$ is the source term. We aim to learn the solution operator of the PDE system, ${\psi}:\mathbb{F}\rightarrow\mathbb{U}$ where $\mathbb{F}$ and $\mathbb{U}$ are two functional spaces, using a training dataset $\mathcal{D}=\{({\bf f}_{n},{\bf u}_{n})\}_{n=1}^{N}$, which includes different instances of $u(\cdot)$ and $f(\cdot)$ sampled/discretized at a set of locations. Once the model is trained, it can be used to directly predict the solution function $u$ for new instances of the input $f$, offering a much more efficient alternative to running numerical solvers from scratch. However, the training dataset still needs to be generated offline using numerical solvers.
Fourier Neural Operator (FNO)  (Li et al., 2020b) is a popular neural network architecture for operator learning, especially in solving PDEs. For a given discretized input function $\bf f$}, FNO first employs a linear layer on each component of $\bf f$} at its respective sampling location, thereby lifting the input into a higher-dimensional channel space. The core of FNO is the Fourier layer, which performs a linear transformation followed by a nonlinear activation within the functional space, $h(\mathbf{x})\leftarrow\sigma\left(\mathcal{W}h(\mathbf{x})+\int\kappa(\mathbf{x}-\mathbf{x}^{\prime})h(\mathbf{x}^{\prime})d\mathbf{x}^{\prime}\right)$, where $h(\mathbf{x})$ is the input to the Fourier layer, $\kappa(\cdot)$ the integration kernel, and $\sigma(\cdot)$ the activation function. The functional convolution is computed using the convolution theorem: $\int\kappa(\mathbf{x}-\mathbf{x}^{\prime})h(\mathbf{x}^{\prime})d\mathbf{x}^{\prime}=\mathcal{F}^{-1}[\mathcal{F}[\kappa]\cdot\mathcal{F}[h]](\mathbf{x})$, where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively. FNO performs Fast Fourier Transform (FFT) on $h$, multiplies it with the discretized kernel in the frequency domain, and then applies the inverse FFT. After multiple Fourier layers, FNO employs another linear layer to project the latent channels to the original space for prediction.
@@
Deep Operator Network (DONet) (Lu et al., 2021) is another prominent work in operator learning. The architecture of a DONet is structured into two components: a branch net and a trunk net, learning representations for the input functions and querying locations, respectively. Consider an input function $f({\bf x})\in\mathbb{F}$ sampled at $m$ locations $\{{\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{m}\}$ and an output function $u\in\mathbb{U}$. The branch net receives the values $f({\bf x}_{1}),f({\bf x}_{2}),\cdots,f({\bf x}_{m})]$ and outputs a feature representation $b_{1},b_{2},\cdots,b_{p}]^{\top}\in\mathbb{R}^{p}$. Concurrently, the trunk network processes a querying location $\bf x$} and outputs another feature vector $t_{1},t_{2},\cdots,t_{p}]^{\top}\in\mathbb{R}^{p}$. The output function value at $\bf x$} is predicted as a sum of products of the corresponding elements from the branch and trunk nets, $\mathcal{G}[f]({\bf x})\approx\sum_{k=1}^{p}b_{k}t_{k}$, where $\mathcal{G}$ is the learned operator mapping $f$ to $u$.
@@
Physics-Informed Neural Operator (PINO) (Wang et al., 2021; Li et al., 2021) has recently emerged as a promising approach to address the data scarcity issue in operator learning. PINO embeds physical laws — typically governing equations — into the learning process. The incorporation of physics not only enhances the model’s adherence to ground-truth phenomena but also reduces its dependency on extensive training data. Mathematically, the integration of physics into the learning process can be viewed as adding a regularization term in the loss function. Let $\mathcal{L}_{\text{data}}$ represent the standard data-fitting loss term (e.g., the mean squared error between the predicted and actual outputs), and the physics-informed term $\mathcal{L}_{\text{physics}}$ be the residual of the governing PDEs evaluated at the neural network’s outputs. The loss function $\mathcal{L}$ for a PINO model is expressed as
@@ 3 Methodology
In the absence of physics knowledge (i.e., PDE system equation 1), it is impossible to construct the physics loss term as in PINO. To address this challenge, we propose a “pseudo” physics-informed operator learning framework that extracts useful physics representation from data so as to enhance operator learning. This framework is motivated by relatively complex applications, where data is often costly and/or limited while the underlying physics is hard to fully understand.
@@ 3.1 Pseudo Physics System Learning
As the first step, we propose a novel approach to learn the physics system using scarce training data. Our key observation is that, although the mapping from $f$ to $u$ can be intricate and may necessitate global information across the entire domain (e.g., in linear PDEs, $u$ is an integration of the Green’s function multiplied with $f$ over the domain), the underlying PDE system equation 1 simplifies to a local combination of $u$ and its derivatives. We therefore design a neural network $\phi$ to approximate the general form of $\mathcal{N}$,
@@
The inherent local combinatorial nature of the PDE representation decouples the values of $u$ and its derivatives across different sampling locations, thereby significantly reducing the learning complexity and the amount of training data required. For instance, consider sampling the input function $f$ and output function $u$ on a $128\times 128$ grid. A single pair of discretized input and output functions, denoted as $({\bf f},{\bf u})$, is typically insufficient for training a neural operator, because learning the mapping $f\rightarrow u$ requires abundant global information. However, this pair can be decomposed into $128\times 128=16,384$ training data points across various (spatial and temporal) locations to train $\phi$ as outlined in equation 3. We use each point of the grid to construct a training sample. This decomposition provides rich information about the local relationships between those derivatives. Therefore, even with a small number of $({\bf f},{\bf u})$ pairs, we hypothesize that the learning of the PDE system $\mathcal{N}$ through our formulation in equation 3 can still yield promising accuracy in predicting $f(\cdot)$ as in equation 1.
@@
We use finite difference to obtain the derivatives of each $u_{n}$, namely, $S_{k}(u_{n})$ ($1\leq k\leq Q$), and then pass these inputs to the network $\phi$ to compute the prediction. Specifically, we employ centered finite differences to approximate the derivatives of $u$, a classical scheme with well-established accuracy and sample complexity. The approximation error scales as $O(h^{2})$ for all derivative orders, where $h$ is the grid spacing. To achieve an error $\leq\epsilon$, the required number of grid points is $n=\Theta(\epsilon^{-d/2})$, where $d$ is the input dimension (Mitchell & Griffiths, 1980). It is worth-noting that while we can separate out training examples across grid points, these examples are not independent to each other. The effective sample size (ESS) for training our $\phi$-network is smaller than the total number of grid points. To compensate numerical errors of the finite difference and leverage inter-dependencies between the training examples, we incorporate a convolution layer to integrate the neighborhood information (see Fig. 1 top). Let $\hat{S}_{k}(u_{n})$ represents the finite difference approximation of $S_{k}(u_{n})$ ($1\leq k\leq Q)$; for $k=0$, we define $\hat{S}_{k}(u_{n})=u_{n}$. Each $\hat{S}_{k}(u_{n})$ is treated as an input channel. After applying the convolution, the output at each sampling location ${\bf x}_{j}$ for channel $k$ is given by
@@
where $\text{nei}(\cdot)$ denotes the neighborhood sampling locations defined in the convolution filter, $\Delta_{ij}$ is the relative distance between ${\bf x}_{i}$ and ${\bf x}_{j}$, and $w_{c}(\Delta_{ij})$ is the corresponding filter weight. Each $\overline{S}_{k}({\bf x}_{j})$ can be viewed as an interpolation, providing a new approximation of $S_{k}({\bf x}_{j})$ by incorporating all neighborhood values of $u_{n}({\bf x}_{j})$ and their finite difference derivative approximations. The convolution results are then passed into subsequent layers. The interpolation (filter) weights are jointly learned from data. This design allows us to integrate additional information into the inputs to facilitate the learning of $\phi$. Empirical results confirm that incorporating this convolution layer improves the prediction accuracy; refer to the ablation study in Section 5.2 and Table 7(a). Next, at each sampling location ${\bf x}_{j}$, we employ fully connected layers (i.e., MLP) to combine all $\{\overline{S}_{k}({\bf x}_{j})\}_{k=0}^{Q}$ in an arbitrarily flexible way to predict $f({\bf x}_{j})$; see Fig. 1 bottom.
As a neural network, we note that the $\phi$-network, has sufficient representation power to approximate any PDE operator of the form
$L(x,u(x),S_{1}(u)(x),\ldots,S_{Q}(u)(x)),$
where $S_{i}$ denotes derivatives, e.g., $L=u_{xx}+uu_{x}+u^{2}-1+\cos(u)$. Because most practical PDEs are smooth analytic functions of $u$ and its derivatives, the universal approximation theorem guarantees that such mappings can be accurately approximated by a feed-forward network of the following form:
$f(x)=\sum_{i=1}^{N}a_{i},\sigma(w_{i}^{\top}x+b_{i}),$
given sufficient neurons and a non-polynomial activation (Cybenko, 1989) (Hornik, 1991). The sample complexity $m(\epsilon,\delta)$ is the smallest number of samples $m$ ensuring, with probability at least $1-\delta$,
$|R(h)-\hat{R}(h)|\leq\epsilon,\quad\forall h\in\mathcal{H},$
where $R(h)$ and $\hat{R}(h)$ denote generalization and training errors, respectively. From computational learning theory (Shalev-Shwartz & Ben-David, 2014),
$m(\epsilon,\delta)=O\left(\frac{d_{VC}+\log(1/\delta)}{\epsilon^{2}}\right),$
where $d_{VC}$ is the VC dimension of the network. For a network with $W$ parameters and piecewise-linear activations (e.g., ReLU), $d_{VC}=O(W\log W)$; for smooth activations, $d_{VC}=O(W^{2})$. Therefore, the learned neural network mapping $\phi:u\rightarrow f$, although black-box in nature, is expressive enough to encapsulate valuable physics knowledge inherent in the data employed for operator learning.
Our method can be readily adapted to other numerical approaches for derivative approximation. For instance, when functions are irregularly sampled, smooth function estimators such as kernel interpolation (Long et al., 2024), the RBF-FD method (Fornberg et al., 2013; Fornberg & Flyer, 2015), or Bayesian B-splines (Sun et al., 2022) can be used to estimate gradient information directly from data. These estimated gradients can then serve as inputs to our PDE neural network $\phi$ for further training.
@@ 3.2 Coupling Neural Operator with Pseudo Physics
Next, we leverage the pseudo physics laws embedded in the learned mapping $\phi:u\rightarrow f$ to enhance neural operator learning. Specifically, we use $\phi$ to reconstruct $f$ from the $u$ predicted by the neural operator. In this way, our approach uses the physics learned in the previous step to incorporate a reconstruction error into the learning of the neural operator parameters; see Fig. 2 for an illustration.
@@
where $N^{\prime}$ is the number of input function samples. To be more specific, the PDE network $\phi$ acts purely as a model-based regularizer, without introducing any additional labeled data. In practice, it is realistic to sample as many as possible new input $f^{\prime}$ from the same input distribution, but it is often not realistic to acquire all the corresponding ground-truth solutions $u^{\prime}$. Thus, we sampled $f^{\prime}$, pass it through the current operator network $\psi$ to obtain a predicted solution $\hat{u}^{\prime}$, and then feed $\hat{u}^{\prime}$ into the PDE network $\phi$ to produce a reconstructed source term $\tilde{f}^{\prime}$. We use the discrepancy $\|\tilde{f}^{\prime}-f^{\prime}\|_{2}$ as a regularization term, as formalized in  equation 5. It simply encourages $\psi$’s predictions to lie on the manifold of solutions that are consistent with the PDE representation $\psi$.
@@
2:Train a preliminary pseudo physics network $\phi$ with loss ${\mathcal{L}}_{\phi}$ in equation 3.
@@
5:  Fix $\phi$, fine tune $\psi$ with loss equation 5.
6:  Fix $\psi$, fine tune $\phi$ with loss equation 5.
7:until maximum iterations are done or convergence.